OpenStudy (anonymous):

Hi need help with three math problems Solve the formula for the given variable 54. 5x-2y=22 for y 53. 4x+3y=12 for y 51. 9x+3y=6 for y

OpenStudy (texaschic101):

54. 5x - 2y = 22 for y first subtract 5x from each side -2y = 22 - 5x then divide everything by -2 to get y by itself y = (22 - 5x)/-2 or y = -11 + 5/2x 53. 4x + 3y = 12 for y subtract 4x from each side 3y = 12 - 4x divide the equation by 3 y = (12 - 4x)/3 or y = 4 - 4/3x 51. 9x + 3y = 6 for y subtract 9x from each side 3y = 6 - 9x divide by 3 y = (6 - 9x)/3 or y = 2 - 3x